# J-VAR: Analysis of RR Lyrae light curves in seven optical bands

### v1

@@ Abstract
Context. RR Lyrae stars are key distance indicators and tracers of old stellar populations in the Milky Way and its satellites. Accurate period and amplitude measurements are essential for constraining stellar evolution and mapping Galactic structure. The Javalambre VARiability (J-VAR) survey provides valuable time-series data across seven optical bands, including grigri and four medium and narrow bands.
Aims. Our goal is to construct and analyze light curves for RR Lyrae stars identified in the first data release (DR1) of the J-VAR survey using the Gaia DR3 variable stars catalog as a reference.
Methods. The light curves of $315$ RR Lyrae were analyzed by fitting templates from the Sloan Digital Sky Survey (SDSS) multiband template library. The periods and amplitudes for the seven bands in J-VAR were independently obtained from the best-fitted templates.
Results. The J-VAR periods show strong agreement with Gaia DR3 values, which confirms the reliability of the applied methodology. The Bailey diagram for each J-VAR filter reveals the expected trend for RRab and RRc stars, with larger pulsation amplitudes at bluer wavelengths. After normalizing by the $r$-band amplitude, an exponential trend is found, with the bluer J-VAR filter centered at $395$ nm having twice the amplitude of the reddest J-VAR passband at $861$ nm. The normalized amplitudes of the RR Lyrae stars from Gaia and the Zwicky Transient Facility (ZTF) are consistent with the J-VAR trend. Finally, the SDSS templates derived from broadbands also provide a proper description for the medium and narrow band light curves.
Conclusions. The J-VAR RR Lyrae catalog offers reliable pulsation parameters and light curves in seven optical filters, allowing the systematic study of amplitude trends from $395$ to $860$ nm and offering a unique benchmark for theoretical models.
@@ 1 Introduction
Variable stars are among the most valuable tools in astronomy. These stars serve as reliable standard candles for determining cosmic distances (Cacciari & Clementini, 2003) and offer critical insights into stellar evolution and Galactic structure (Catelan, 2009; Catelan & Smith, 2015). As excellent tracers of old stellar populations, variable stars such as RR Lyrae, located in the halo of the Milky Way, are ideal for studying their formation and evolution (Pietrukowicz, 2016; Prudil et al., 2022). Furthermore, RR Lyrae stars having periods of pulsations in fundamental frequency are classified as RRab subtype; those having periods in the first overtones are classified as RRc subtype, and those having periods in both frequencies are classified as RRd subtype stars.
Recent large-scale surveys have revolutionized our understanding of RR Lyrae stars. The Gaia (Gaia Collaboration 2016), particularly the third data release (DR3; Gaia Collaboration 2022a), provide precise astrometric, photometric, and spectroscopic data for millions of stars, including RR Lyrae. These massive data have since served as a foundation for several follow-up studies, including Dark Energy Survey Instrument (DESI) spectroscopy (Medina et al., 2025), among others. In parallel, time-domain surveys such as the Zwicky Transient Facility (ZTF; Bellm et al. 2019) have provided optical photometry over large sky areas, enabling the detection and detailed light curve analysis of variable stars, including RR Lyrae, on short timescales. The ZTF camera has a wide field of view ($\sim 47$ deg2; Dekany et al. 2020) that captures the pulsation properties and phase-resolved features of these stars.
RR Lyrae stars also exhibit a well-known decrease in pulsation amplitude with increasing wavelength, a trend arising from the reduced contrast between maximum and minimum light at longer wavelengths (Catelan, 2009). Most of the photometric studies characterizing this effect have been conducted using broadband filters, which, while efficient for large-scale surveys, blend information from different spectral features. Notable exceptions include narrow-band studies in the Strömgren and Caby photometric systems (Baird, 1996; Peña et al., 2009, 2012), which have demonstrated the diagnostic potential of narrower passbands in probing temperature-sensitive and line-specific variations across the pulsation cycle.
@@
This work analyses the RR Lyrae stars identified in J-VAR DR1 confirmed through cross-matching with the Gaia DR3 variable star catalog. This ensures a robust classification and provides the periods from the Gaia DR3 along with the epochs of maximum brightness. Later, the magnitudes from J-VAR DR1 were fitted using the templates from the Sloan Digital Sky Survey (SDSS; York et al. 2000) multiband template library (Sesar et al., 2010). The J-VAR periods and amplitudes are used to study the variation of the RR Lyrae light curve properties along the optical range and to our best knowledge for the first time in the $J0515$, $J0660$ and $J0861$ passbands; being the $J0395$ similar to the $Ca$ filter studied by Baird (1996).
This paper is organized as follows. The J-VAR DR1 data are presented in Sect. 2, the methodology is described in Sect. 3, and the results are detailed in Sect. 4. Furthermore, the discussions and conclusions are in Sects.5 and 6, respectively. Finally, Appendix A describes the catalog for light curve analysis of the RR Lyrae stars in J-VAR DR1. All the magnitudes are in the AB scale (Oke & Gunn, 1983).
@@ 2 Data
Table 1: J-VAR photometric system
@@ 2.1 J-VAR DR1 Data
J-VAR is conducted from the Observatorio Astrofísico de Javalambre (OAJ, Teruel, Spain; Cenarro et al. 2014) with the 83 cm Javalambre Auxiliary Survey Telescope (JAST80) and T80Cam, a wide-field camera of 9.2k × 9.2k pixels covering a $2$ deg2 field of view with a pixel scale of 0.550.55″pix-1 (Marín-Franch et al., 2015). J-VAR delivers time-series photometry in seven optical bands across 33 epochs in fields already observed by the single-epoch Javalambre Photometric Local Universe Survey (J-PLUS; Cenarro et al. 2019). The observations were taken on non-photometric nights, and the inclusion of four narrow bands enables detailed light curve modeling and spectral energy distribution (SED; Pickles 1998, Nemec et al. 2013, Muraveva et al. 2018) covering spectral features not typically available in other large-scale variability surveys.
J-VAR DR1111https://archive.cefca.es/catalogues/jvar-dr1 includes light curves for 1,335,279 objects classified as stars or quasars by BANNJOS (Bayesian Artificial Neural Networks for the Javalambre Observatory Surveys; del Pino et al. 2024) in J-PLUS, as detailed in Pyrzas et al., in prep. The data provide a signal-to-noise ratio greater than $10$ for objects brighter than $19$ mag in the $r$-band.
Starting from the full J-VAR DR1 catalog, some quality cuts were applied. Unreliable photometry estimation is filtered out by excluding magnitudes exceeding $25$ mag in any filter. Measurements with relative errors greater than 2020% of the magnitude or with absolute errors larger than $0.5$ mag were also discarded. Additionally, custom root mean square (RMS) thresholds (Pyrzas et al., in prep.) were applied to each band to remove outliers.
@@ 2.2 Gaia DR3 Variable Star Catalog
Gaia is a space observatory dedicated to mapping over 1.8 billion stars in the Milky Way. It monitors stellar positions, motions, and brightness variations, enabling large-scale studies of variable stars (Gaia Collaboration et al., 2022b).
The Gaia DR3 variable star catalog $I/358/vrrlyr$ (Gaia Collaboration et al., 2023) has $271,779$ variable stars identified as RR Lyrae. It comprises $175,350$ RRab, $94,422$ RRc, and $2,007$ RRd subtypes. The catalog provides information for photometry in Gaia $G$-band, periods in fundamental frequency and first overtones, and epochs of maximum light.
2.3 Data Cross-matching and Parameter Estimation
@@
To construct light curves, the Modified Julian Dates (MJD) in J-VAR DR1 were converted to Barycentric Julian Dates (BJD) using functions from the Astropy (Astropy Collaboration et al., 2013a, 2018, 2022), which account for the Earth’s motion relative to the barycenter of the Solar System. The transformation incorporated the coordinates of the observation site, latitude $40.0451^{\circ}$, longitude $1.0005^{\circ}$, and elevation $1957\leavevmode\nobreak\ \text{m}$ using the EarthLocation module in Astropy to ensure precise timing corrections.
Figure 1 shows the Bailey Diagram (Bailey, 1981), which plots Gaia periods versus the Gaia amplitudes, for the entire Gaia DR3 catalog and the $315$ RR Lyrae in the J-VAR DR1 catalog. The amplitudes were assumed as halves of the peak-to-peak magnitudes in the Gaia $G$-band. The alignment of objects in J-VAR DR1 with the sources in the Gaia DR3 variable star catalog is confirmed, providing a representative subset of the initial Gaia sample.
@@ 2.4 SDSS Multiband Template Library
The SDSS is a wide-field imaging and spectroscopic survey that has significantly advanced our understanding of the Milky Way and the distant universe. The SDSS multiband template library for RR Lyrae stars, as presented in Sesar et al. (2010), provides a set of well-defined empirical light curve templates based on observations of $483$ RR Lyrae stars ($379$ RRab and $104$ RRc stars) from SDSS Stripe 82. These templates were constructed using a large sample of well-sampled multiband light curves in the ugrizugriz photometric system, with a median of $30$ observations per band for objects with $g<21$ mag. The template families found in that work show a variety of shapes of the light curves. The study demonstrated that RR Lyrae stars, particularly RRab and RRc subtypes, exhibit characteristic light curve shapes that can be effectively modeled using these templates. The library improves period determination, classification accuracy, and distance estimates for RR Lyrae stars, aiding studies of the Galactic structure, especially in the halo. The SDSS template library was used to fit the J-VAR DR1 multiband light curves of each RR Lyrae star.
@@ 3.1 Phase-folded Light Curves
Following the cross-match between the J-VAR DR1 data and the Gaia DR3 variable star catalog, a combined catalog was created. However, this initial version did not include phase-folded light curves. To address this, an additional step was implemented using a custom pipeline that phase-folds the light curves from phase 010-1 using a reference period, normalized over the pulsation cycle. Each light curve is plotted twice (020-2 in phase) to display the variability pattern. Figure 2 shows five representative examples generated using this method and the Gaia periods for reference.
@@ 3.2 Parameter Estimation
The light curves of the $315$ RR Lyrae stars in the J-VAR DR1 + Gaia DR3 catalog were analyzed using a robust template-fitting pipeline. The workflow integrates widely used tools: Astropy (Astropy Collaboration et al., 2013b, 2018, 2022) for time and coordinate transformations, Gatspy (Vanderplas, 2015) for period determination and multiband light curve modeling, NumPy (Harris & et al., 2020) for numerical calculations, and astroML (VanderPlas et al., 2012) for statistical methods, including period refinement. Several Gatspy modules were modified to accelerate period determination through parallelization and to enable the retrieval of the best-fitting template label.
Table 2 outlines the band configurations of J-VAR DR1 and SDSS used for the template fitting. The broad bands grigri were aligned with their direct SDSS counterparts, while the medium and narrow bands were mapped to the closest SDSS filters by effective wavelength. The script fits each light curve over a period range from 010-1 days using the RRLyraeTemplateModeler. Figure 3 shows the light curve analysis for two representative RR Lyrae stars in the catalog.
The template-fitting process provided independent parameter estimates across all seven optical bands. These include the identifier of the best-fitting template, periods of the best-fitted templates (also known as J-VAR periods), and the corresponding $\chi^{2}$ value. As part of the catalog, J-VAR amplitudes, defined as half of the peak-to-peak magnitudes, were also recorded.
@@ 4 Results
Figure 5: Normalized amplitudes vs. wavelength. The points represent the medians of the J-VAR amplitudes for each band normalized by that in the $r$-band. The figure also shows the Gaia $G$-band and the ZTF $g$ and $r$ bands for comparison and were not included in the fit. The dashed black line indicates the exponential fit. Error bars correspond to the 25th ($p_{25}$) and 75th ($p_{75}$) percentiles. N denotes the number of valid amplitude ratios computed per band.
@@ 4.1 Bailey diagram
The Bailey diagram with J-VAR amplitudes and J-VAR periods (Figure 4) reveals a clear separation between RRab and RRc subtypes across both the broad (grigri) and narrow bands. Moreover, the characteristic locus for RRab stars is more distinctly visible compared to that for RRc stars. The J-VAR amplitudes in the bluer passbands are larger than in the redder ones, both for the broad and the medium-to-narrow filters, following the known trend with wavelength (e.g. Catelan, 2009). We recall that each filter was analyzed independently.
@@ 4.2 Normalized amplitudes across wavelengths
To quantitatively study the observed trend of decreasing amplitude with increasing wavelength, the normalized amplitude with respect to the $r$- band amplitude was computed, noted $\overline{A}$. This permits removing the main trends present in the Bailey Diagrams and enhances relative differences across passbands. The interquartile ranges (IQR; defined as the difference between the 75th and 25th percentiles, $p_{75}-p_{25}$) of the normalized amplitudes, along with their medians, are summarized in Table 3 for all seven optical bands in J-VAR DR1. A clear downward trend was observed from short to long wavelengths, indicating that normalized amplitude decreases as wavelength increases, suggesting that the fluxes in bluer bands are more sensitive to temperature variations than those in the redder bands.
@@
where $\overline{A}(\lambda)$ is the normalized amplitude and $\lambda$ is the wavelength in nanometers. This relation captures the expected physical behavior of RR Lyrae stars, where variability amplitude diminishes at longer wavelengths due to the lower temperature sensitivity of the flux described by the Planck Function (Planck, 1901).

### v2

@@ Abstract
Context. RR Lyrae stars, with their accurate period and amplitude measurements, constrain stellar evolution and map Galactic structure. The Javalambre VARiability (J-VAR) survey is the time-domain extension of the Javalambre Photometric Local Universe survey, which provides time-series data across seven optical bands, including grigri and four medium and narrow bands.
Aims. Our goal is to construct and analyze light curves for RR Lyrae stars identified in the J-VAR’s first data release using the Gaia third data release (DR3) Variable Stars catalog as a reference.
Methods. The light curves of $315$ RR Lyrae were analyzed by fitting templates from the Sloan Digital Sky Survey Multiband Template Library. The periods and amplitudes for the seven bands in J-VAR were independently obtained from the best-fitted templates.
Results. The J-VAR periods show strong agreement with Gaia DR3 values. The Bailey diagram for each J-VAR filter shows larger pulsation amplitudes at bluer wavelengths. Amplitudes, after normalizing by the $r$-band amplitude, show an exponential trend, with the bluer J-VAR filter centered at $395\penalty 10000\ \text{nm}$ having twice the amplitude of the reddest J-VAR passband at $861\penalty 10000\ \text{nm}$. The normalized amplitudes of the RR Lyrae stars from Gaia and the Zwicky Transient Facility are consistent with the J-VAR trend. Finally, the SDSS templates derived from broadbands also provide a proper description for the medium and narrow band light curves.
Conclusions. The J-VAR RR Lyrae catalog offers reliable pulsation parameters and light curves in seven optical filters, allowing the systematic study of amplitude trends from $395\penalty 10000\ \text{nm}$ to $860\penalty 10000\ \text{nm}$.
@@ 1 Introduction
Variable stars are among the most valuable tools in astronomy. These stars serve as reliable standard candles for determining cosmic distances (CacciariClementini2003) and offer critical insights into stellar evolution and Galactic structure (Catelan2009; catelan2015). As excellent tracers of old stellar populations, variable stars such as RR Lyrae, located in the halo of the Milky Way, are ideal for studying their formation and evolution (Pietrukowicz2016; Prudil2022). Furthermore, RR Lyrae stars having pulsation periods in fundamental frequency are classified as RRab subtype; those having periods in the first overtones are classified as RRc subtype, and those having periods in both frequencies are classified as RRd subtype stars.
Recent large-scale surveys have revolutionized our understanding of RR Lyrae stars. The Gaia (Gaia Collaboration GaiaMission16), particularly the third data release (DR3; Gaia Collaboration GaiaDR3_2022), provide precise astrometric, photometric, and spectroscopic data for millions of stars, including RR Lyrae. These massive data have since served as a foundation for several follow-up studies, including Dark Energy Spectroscopic Instrument (DESI) (Medina2025_DESI_Y1_RRL), among others. In parallel, time-domain surveys such as the Zwicky Transient Facility (ZTF; Bellm2019) have provided optical photometry over large sky areas, enabling the detection and detailed light curve analysis of variable stars, including RR Lyrae, on short timescales.
RR Lyrae stars also exhibit a well-known decrease in pulsation amplitude with increasing wavelength, a trend arising from the reduced contrast between maximum and minimum light at longer wavelengths (Catelan2009). Most of the photometric studies characterizing this effect have been conducted using broadband filters, which, while efficient for large-scale surveys, blend information from different spectral features. Notable exceptions include narrow-band studies in the Strömgren and Caby photometric systems (baird96; pena09; pena12), which have demonstrated the diagnostic potential of narrower passbands in probing temperature-sensitive and line-specific variations across the pulsation cycle.
@@
This work analyses the RR Lyrae stars identified in J-VAR DR1 confirmed through cross-matching with the Gaia DR3 Variable Star catalog. This ensures a robust classification and provides the periods from the Gaia DR3 along with the epochs of maximum brightness. The magnitudes from J-VAR DR1 were fitted using the templates from the Sloan Digital Sky Survey (SDSS; york2000) multiband template library (Sesar2010). The J-VAR periods and amplitudes are used to study the variation of the RR Lyrae light curve properties along the optical range and to our best knowledge for the first time in the $J0515$, $J0660$ and $J0861$ passbands; being the $J0395$ similar to the $Ca$ filter studied by baird96.
This paper is organized as follows. The J-VAR DR1 data, Gaia Variable Star catalog, SDSS Multiband Template Library, and ZTF data are presented in Sect. 2, the methodology is described in Sect. 3, and the results are detailed in Sect. 4. Furthermore, the discussions and conclusions are in Sects.5 and 6, respectively. Finally, Appendix A describes the catalog for light curve analysis of the RR Lyrae stars in J-VAR DR1. The magnitudes presented throughout this work are in the AB scale (oke83).
@@ 2 Data
Table 1: J-PLUS Photometric Filter Set Adopted by J-VAR
@@ 2.1 J-VAR DR1 Data
J-VAR is conducted from the Observatorio Astrofísico de Javalambre (OAJ, Teruel, Spain; oaj) with the 83 cm Javalambre Auxiliary Survey Telescope (JAST80) and T80Cam, a wide-field camera of 9.2k × 9.2k pixels covering a $2$ deg2 field of view with a pixel scale of 0.550.55″pix-1 (t80cam). J-VAR delivers time-series photometry in seven optical bands across 33 epochs in fields already observed by the single-epoch Javalambre Photometric Local Universe Survey (J-PLUS; Cenarro2019). The observations were taken on non-photometric nights, and the inclusion of four narrow bands enables detailed light curve modeling and spectral energy distribution (SED; Pickles1998, Nemec2013, Muraveva2018) that cover spectral features not typically available in other large-scale variability surveys.
J-VAR DR1111https://archive.cefca.es/catalogues/jvar-dr1 includes light curves for 1,335,279 objects classified as stars or quasars by BANNJOS (Bayesian Artificial Neural Networks for the Javalambre Observatory Surveys; bannjos) in J-PLUS, as detailed in Pyrzas et al., in prep. The data provide a signal-to-noise ratio greater than $10$ for objects brighter than $19$ mag in the $r$-band.
Starting from the J-VAR DR1 jvar.Light_curves table (Ederoclite_2025) available at CEFCA’s catalog portal [^2] https://archive.cefca.es/catalogues/jvar-dr1, some quality cuts were applied. Unreliable photometry estimation is filtered out by excluding magnitudes exceeding $25$ mag in any filter. Measurements with relative errors greater than 2020% of the magnitude or with absolute errors larger than $0.5$ mag were also discarded. Additionally, custom root-mean-square thresholds (Pyrzas et al., in prep.) were applied to each band to remove outliers.
@@ 2.2 Gaia DR3 Variable Star Catalog
Gaia is a space observatory dedicated to mapping over 1.8 billion stars in the Milky Way. It monitors stellar positions, motions, and brightness variations, enabling large-scale studies of variable stars (GaiaCollaboration2022).
The Gaia DR3 Variable Star catalog $I/358/vrrlyr$ (vrrlyr_catalog_2023) has $271,779$ variable stars identified as RR Lyrae. It comprises $175,350$ RRab, $94,422$ RRc, and $2,007$ RRd subtypes. The catalog provides information for photometry in Gaia $G$-band, periods in fundamental frequency and first overtones, and epochs of maximum light.
2.3 Data Cross-matching
@@
To construct light curves, the Modified Julian Dates (MJD) in J-VAR DR1 were converted to Barycentric Julian Dates (BJD) using functions from Astropy (AstropyCollaboration2013; astropy:2018; astropy:2022), which account for the Earth’s motion relative to the barycenter of the Solar System. The transformation incorporated the coordinates of the observation site, latitude $40.0451^{\circ}$, longitude $1.0005^{\circ}$, and elevation $1957\penalty 10000\ \text{m}$ using the EarthLocation module in Astropy to ensure precise timing corrections.
Fig. 1 shows the Bailey Diagram (Bailey1981), which plots Gaia periods versus the Gaia amplitudes, for the entire Gaia DR3 catalog and the $315$ RR Lyrae in the J-VAR DR1 catalog. The amplitudes were assumed as halves of the peak-to-peak magnitudes in the Gaia $G$-band. The alignment of objects in J-VAR DR1 with the sources in the Gaia DR3 Variable Star catalog is confirmed, providing a representative subset of the initial Gaia sample. The gap between the RRab and RRc populations may be partly influenced by residual systematics in the Gaia-J-VAR amplitude calibration (Clementini_2023), or it may simply reflect the intrinsic Oosterhoff dichotomy (Oosterhoff_1939, catelan2015), which naturally produces distinct RRab and RRc loci.
@@ 2.4 SDSS Multiband Template Library
The SDSS is a wide-field imaging and spectroscopic survey that has significantly advanced our understanding of the Milky Way and the distant universe. The SDSS Multiband Template Library for RR Lyrae stars, as presented in Sesar2010, provides a set of well-defined empirical light curve templates based on observations of $483$ RR Lyrae stars ($379$ RRab and $104$ RRc stars) from SDSS Stripe 82. These templates were constructed using a large sample of well-sampled multiband light curves in the ugrizugriz photometric system, with a median of $30$ observations per band for objects with $g<21$ mag. The template families found in that work show a variety of shapes of the light curves. The study demonstrated that RR Lyrae stars, particularly RRab and RRc subtypes, exhibit characteristic light curve shapes that can be effectively modeled using these templates. The library improves period determination, classification accuracy, and distance estimates for RR Lyrae stars, aiding studies of the Galactic structure, especially in the halo. The SDSS template library was used to fit the J-VAR DR1 multiband light curves of each RR Lyrae star.
2.5 The ZTF Catalog
To complement the J-VAR and Gaia photometry, the light-curve information from the Zwicky Transient Facility (ZTF; Masci2019) was also used. ZTF provides multi-epoch photometry in the $g$ and $r$ bands, with passband profiles broadly similar to the corresponding J-VAR filters. Amplitude measurements for RR Lyrae stars were taken from the ZTF DR3 catalog (J/ApJS/249/18) [^3] https://vizier.cds.unistra.fr/viz-bin/VizieR-3?-source=J/ApJS/272/31 (Chen2020_ZTF_Catalog). These values are used for cross-survey comparisons in Sect. 5.2.
@@ 3.1 Phase-folded Light Curves
Following the cross-match between the J-VAR DR1 data and the Gaia DR3 Variable Star catalog, a combined catalog was created. However, this initial version did not include phase-folded light curves. To address this, an additional step was implemented using a custom pipeline that phase-folds the light curves from phase 0 to phase $1$ using a reference period, normalized over the pulsation cycle. Each light curve is plotted twice (020-2 in phase) to display the variability pattern. Fig. 2 shows five representative examples generated using this method and the Gaia periods for reference.
@@ 3.2 Parameter Estimation
The light curves of the $315$ RR Lyrae stars in the J-VAR DR1 + Gaia DR3 catalog were analyzed using a robust template-fitting pipeline. The workflow integrates widely used tools: Astropy for time and coordinate transformations, Gatspy (Vanderplas2015) for period determination and multiband light curve modeling, NumPy (Harris_2020) for numerical calculations, and astroML (Vanderplas2012) for statistical methods, including period refinement. Several Gatspy modules were modified to accelerate period determination through parallelization and to enable the retrieval of the best-fitting template label.
Tab. 2 outlines the band configurations of J-VAR DR1 and SDSS used for the template fitting. The broad bands grigri were aligned with their direct SDSS counterparts, while the medium and narrow bands were mapped to the closest SDSS filters by effective wavelength. The script fits each light curve over a period range from 010-1 days using the RRLyraeTemplateModeler. Fig. 3 shows the light curve analysis for two representative RR Lyrae stars in the catalog.
The template-fitting process provided independent parameter estimates across all seven optical bands. These include the identifier of the best-fitting template, periods of the best-fitted templates (also known as J-VAR periods), and the corresponding $\chi^{2}$ value, which is reported for completeness but not further used in this study. The uncertainties on the J-VAR periods are extremely small ($\lesssim 10^{-6}$$10^{-5}$ days), and are therefore negligible for this analysis. As a part of the catalog, J-VAR amplitudes, defined as half of the peak-to-peak magnitudes, were also recorded.
@@ 4 Results
Figure 5: Normalized amplitudes vs. wavelength. The points represent the medians of the J-VAR amplitudes for each band normalized by that in the $r$-band. The figure also shows the Gaia $G$-band and the ZTF $g$ and $r$ bands (normalized by the J-VAR $r$-band) for comparison and were not included in the fit. The dashed black line indicates the exponential fit. Error bars correspond to the 25th ($p_{25}$) and 75th ($p_{75}$) percentiles. N denotes the number of valid amplitude ratios computed per band.
@@ 4.1 Bailey diagram
The Bailey diagram with J-VAR amplitudes and J-VAR periods (Figure 4) reveals a clear separation between RRab and RRc subtypes across both the broad (grigri) and narrow bands. Moreover, the characteristic locus for RRab stars is more distinctly visible compared to that for RRc stars. The J-VAR amplitudes in the bluer passbands are larger than in the redder ones, both for the broad and the medium-to-narrow filters, following the known trend with wavelength (e.g. Catelan2009). We recall that each filter was analyzed independently.
@@ 4.2 Normalized amplitudes across wavelengths
To quantitatively study the observed trend of decreasing amplitude with increasing wavelength, the normalized amplitude with respect to the $r$- band amplitude was computed, noted $\overline{A}$ (see Fig. 5). This permits removing the main trends present in the Bailey Diagrams and enhances relative differences across passbands. The interquartile ranges (IQR; defined as the difference between the 75th and 25th percentiles, $p_{75}-p_{25}$) of the normalized amplitudes, along with their medians, are summarized in Tab. 3 for all seven optical bands in J-VAR DR1. A clear downward trend was observed for increasing wavelengths, indicating that normalized amplitude decreases as wavelength increases, suggesting that the fluxes in bluer bands are more sensitive to temperature variations than those in the redder bands.
@@
where $\overline{A}(\lambda)$ is the normalized amplitude and $\lambda$ is the wavelength in nanometers. This relation captures the expected physical behavior of RR Lyrae stars, where variability amplitude diminishes at longer wavelengths due to the lower temperature sensitivity of the flux described by the Planck Function (planck1901).